# Learning stochastic dynamics from snapshots through regularized unbalanced optimal transport

### v1

@@ 5 Learning RUOT through Deep Learning
$=\mathbb{E}_{\boldsymbol{x_{0}}\sim p_{0}}\int_{0}^{1}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]e^{\int_{0}^{t}g(\boldsymbol{x},s)\mathrm{d}s}\mathrm{d}t,$
which we compute through Monte Carlo sampling and Neural ODE solver.
5.2 Reconstruction Loss
The reconstruction loss aims to match the final distribution in Theorem 3.2, which is composed of two parts: the mass matching loss and the optimal transport loss in matching the normalized distributions. Specifically, we first address the mass matching loss to align the total number of cells. Subsequently, we normalize the distributions according to the matched masses, and then use these weights to perform optimal transport matching. This process is similar to using unbalanced optimal transport to evaluate the distance between two distributions.
For the first part, we propose a local mass-matching strategy. We denote $A_{i}$ ($i=0,1,\cdots T-1$) as the dataset observed at different time points, and define $\phi_{\theta}:\mathbb{R}^{d}\times\mathcal{T}\rightarrow\mathbb{R}^{d|\mathcal{T}|}$ as the function representing the Neural ODE, with $\mathcal{T}$ being a set of time indices. Starting with an initial set $A_{0}$, the Neural ODE function $\phi_{\theta}$ predicts the subsequent sets of data points over the time indices in $\mathcal{T}$. Specifically, the predicted sets $\widehat{A}_{1},\ldots,\widehat{A}_{T-1}$ are obtained by applying the Neural ODE function to $A_{0}$ and the full range of time indices from $1$ to $T-1$, i.e., $\widehat{A}_{1},\ldots,\widehat{A}_{T-1}=\phi_{\theta}\left(A_{0},\{1,\ldots,T-1\}\right).$ Assume the sample size of $N_{0}$ at the initial time point, with the relative total masses at subsequent time points denoted as $n_{t_{k}}=N_{k}/N_{0}$ for $k=0,1,\cdots,T-1$. Along the trajectory, each sampled particle has a weight $w_{i}(t)$ and it follows that
$\frac{\mathrm{d}\log w_{i}(t)}{\mathrm{d}t}=g(\boldsymbol{x}_{i}(t),t),\ \ w_{i}(0)=1/N_{0}.$
We then establish a mapping $h_{k}$ from points at time $t_{k}$ to the sampled particles (total number $N_{0}$) predicted at time $t_{k}$, which links the real data points in $A_{k}$ to its closest point in sampled particles at $\widehat{A}_{k}$, i.e.,
$h_{k}:A_{k}\rightarrow\widehat{A}_{k},\ h_{k}(\boldsymbol{{x}}_{t_{k}})=\text{argmin}_{\boldsymbol{y}\in\widehat{A}_{k}}\|\boldsymbol{{x}}_{t_{k}}-\boldsymbol{y}\|_{2}^{2}.$
The mass matching error at time $t_{k}$ is then defined as
$M_{k}=\sum_{i=1}^{N_{0}}\Big{\|}w_{i}(t_{k})-\text{card}\left(h_{k}^{-1}(\boldsymbol{x}_{i}(t_{k}))\right)\frac{1}{N_{0}}\Big{\|}_{2}^{2}.$
Here $\text{card}(A)$ denotes the cardinality (i.e., the number of elements for finite set) of set $A$. This error metric ensures the matching of local masses. Furthermore, when $M_{t}=0$, we have
$\sum_{i=1}^{N_{0}}w_{i}(t_{k})=\sum_{i=1}^{N_{0}}\text{card}\left(h_{t}^{-1}(\boldsymbol{x}_{i}(t_{k}))\right)/N_{0}=N_{k}/N_{0}=n_{t_{k}}.$
[…]
Once we have determined the weights $w_{i}(t)$, we utilize these weights to perform optimal transport matching of the distributions
$D_{0,T}:=\sum_{k=1}^{T-1}\mathcal{W}_{2}\big{(}\hat{\boldsymbol{w}}^{k},\boldsymbol{w}(t_{k})\big{)},$
where $\hat{\boldsymbol{w}}^{k}=(1/N_{k},1/N_{k},...,1/N_{k})$ is the uniform distribution of $A_{k}$ at time $t$, $\boldsymbol{w}(t_{k})=(w_{1}(t_{k}),w_{2}(t_{k}),...,w_{N_{0}}(t_{k}))/\sum_{i=1}^{N_{0}}w_{i}(t_{k})$ is the predicted weight distribution of sample particles at time $t_{k}$ and $\mathcal{W}_{2}$ represents the Wasserstein distance between $A_{k}$ and $\widehat{A}_{k}$ with normalized distribution defined by $\hat{\boldsymbol{w}}^{k}$ and $\boldsymbol{w}(t_{k})$. Thus the reconstruction loss is
$R_{d}=\lambda_{m}\sum_{k=1}^{T-1}M_{k}+\lambda_{d}D_{0,T},$ (5.20)
where $\lambda_{m}$ and $\lambda_{d}$ are hyperparameters.
5.3 Learning Probability Flow ODE and Score Function
We now integrate all components to derive the deep learning-based solver for learning the probability flow ODE. We utilize fully connected neural networks $\boldsymbol{v}_{\theta},g_{\theta},s_{\theta}$ to estimate $\boldsymbol{v}(\boldsymbol{x},t)$, $g(\boldsymbol{x},t)$ and $\frac{\sigma^{2}}{2}\log{p}(\boldsymbol{x},t)$ respectively. The energy function is computed using Eq. (5.19) and the reconstruction error is computed using Eq. (5.20). Additionally, it is necessary to incorporate a physics-informed loss (PINN-loss) (Raissi et al., 2019) to constrain the relationships among the three neural networks, i.e. Eq. 3.10. Also we utilize a Gaussian mixture model to estimate the initial distribution, ensuring that it satisfies the initial boundary conditions.
$W_{\text{PINN}}=\left\|\partial_{t}p+\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)-g(\boldsymbol{x},t)p\right\|_{2}+\lambda_{w}\left\|p(x,0)-p_{0}\right\|_{2}.$
So overall, the loss function is defined as the weighted sum of the cost, the reconstruction error and the PINN loss:
$L=R_{d}+\lambda_{m}W_{0,T}+\lambda_{p}W_{\text{PINN}},$
with the loss weights hyper-parameters $\lambda_{m}$ and $\lambda_{p}$. Here, $\boldsymbol{x}=\boldsymbol{x}(t)$ is defined by the following equation.
$\begin{cases}\frac{\mathrm{d}\boldsymbol{x}(t)}{\mathrm{d}t}=\boldsymbol{v}(\boldsymbol{x}(t),t),\\ \boldsymbol{x}(t_{i})=\boldsymbol{x}_{i}.\end{cases}$
This allows us to develop a neural network algorithm for solving the RUOT problem (Algorithm 1).
Algorithm 1 Learning Regularized Unbalanced Optimal Transport
[…]
$\widehat{A}_{t+1}\leftarrow\boldsymbol{\boldsymbol{v_{\theta}}}\left(\widehat{A}_{t},t+1\right)$
2:$L_{d}\leftarrow L$
[…]
5.4 Training
Here we adopt a multi-stage training approach to deal with multiple loss terms and stabilize the training process. Initially, we used reconstruction error to train $\boldsymbol{v}$ and $g$, ensuring a required matching. Subsequently, we fixed $\boldsymbol{v}$ and $g$ and employed a flow-matching method to learn an initial, well-optimized log density function $s_{\theta}(\boldsymbol{x},t)$. Following this, we refined the log density function based on the PINN loss. Finally, using $\boldsymbol{v}$, $g$, and the optimized log density function as the starting point, we use Algorithm 1 to obtain the final result.
Our approach to learning the initial log density function is conducted by conditional flow matching (Lipman et al., 2023; Tong et al., 2024a). Previously, we have learnt a matching. Then we perform the dynamics building on this to generate samples, which is akin to conducting an unbalanced optimal transport. We perform conditional flow matching (CFM) within such samples same as (Tong et al., 2024b). Specifically, we choose pair samples $(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ from the optimal transport plan $q(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ and then construct Brownian bridges between these pair samples. We similar first consider $\sigma(t)=\sigma$ is constant. The log density function is then matched with these Brownian bridges, i.e., $p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))=\mathcal{N}\left(\boldsymbol{x};t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0},\sigma^{2}t(1-t)\right)$ and $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))=\frac{t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0}-\boldsymbol{x}}{\sigma^{2}t(1-t)},~{}t\in[0,1].$ We then utilize the fully connected neural networks $s_{\theta}(\boldsymbol{x},t)$ to approximate the log density function $\frac{1}{2}\sigma^{2}\log{p}(\boldsymbol{x},t)$ with a weighting function $\lambda_{s}$. Then we have
$L_{us}=\lambda_{s}^{2}\|\nabla_{\boldsymbol{x}}s_{\theta}-\frac{1}{2}\sigma^{2}\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t)\|_{2}^{2}.$
And the correspongding CFM loss is
$L_{s}=\mathbb{E}_{Q^{\prime}}\lambda_{s}^{2}\|\nabla_{\boldsymbol{x}}s_{\theta}-\frac{1}{2}\sigma^{2}\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))\|_{2}^{2},$
Then we have
$L_{S}=$ $\lambda_{s}(t)^{2}\left\|\nabla_{\boldsymbol{x}}s_{\theta}(\boldsymbol{x},t)-\frac{\sigma^{2}}{2}\nabla_{\boldsymbol{x}}\log p\left(\boldsymbol{x}\mid\boldsymbol{x_{0}},\boldsymbol{x_{1}}\right)\right\|^{2},$
$=\left\|\lambda_{s}(t)\nabla_{\boldsymbol{x}}s_{\theta}(\boldsymbol{x},t)-\lambda_{s}(t)\frac{\sigma^{2}}{2}\nabla_{\boldsymbol{x}}\log p\left(\boldsymbol{x}\mid\boldsymbol{x_{0}},\boldsymbol{x_{1}}\right)\right\|^{2},$
$=\left\|\lambda_{s}(t)\nabla_{\boldsymbol{x}}s_{\theta}(\boldsymbol{x},t)+\epsilon_{1}\right\|^{2},$
where $\epsilon_{1}\sim\mathcal{N}(0,1)$. This is also numerically stable. For the case where $\sigma(t)$ is not constant, a similar approach can be applied. One may refer to their paper for further details.
In summary, by integrating all the discussed methodologies, we derive the Algorithm 2 for training the regularized unbalanced optimal transport in practice.
Algorithm 2 Training Regularized Unbalanced Optimal Transport
[…]
$\widehat{A}_{t+1}\leftarrow\boldsymbol{\boldsymbol{v_{\theta}}}\left(\widehat{A}_{t},t+1\right)$
2:$L\leftarrow R_{d}$
[…]
5:$(\boldsymbol{x}_{0},\boldsymbol{x}_{1})\sim q(\boldsymbol{x}_{0},\boldsymbol{x}_{1});\quad t\sim\mathcal{U}(0,T);\quad\boldsymbol{x}\sim p(\boldsymbol{x},t\mid(\boldsymbol{x}_{0},\boldsymbol{x}_{1}))$ at generated datasets $\widehat{A}_{0},\cdots,\widehat{A}_{T-1}$.
6:$L\leftarrow L_{s}$
7:update $s_{\theta}$ with gradient descent w.r.t. the loss $L$EndFor
8:Conduct Algorithm 1
[…]
In practical implementation, we employ Algorithm 2. The temporal integral and ODEs were numerically solved using the Nerual ODE solver (Chen et al., 2018). The gradients of the loss function with respect to the parameters in the neural networks for $\boldsymbol{v}(\boldsymbol{x},t)$, $g(\boldsymbol{x},t)$ and $s(\boldsymbol{x},t)$ were computed using neural ODEs with a memory-efficient implementation. The Adam optimizer was employed to update the gradients. For computing the Wasserstein distance between discrete distributions, we utilize the implementation provided by the Python Optimal Transport library (POT) (Flamary et al., 2021).
6 Numerical Results
In this section, we evaluate the effectiveness of the unbalanced dynamic diffusion Schrödinger bridge solver on synthetic dynamics from two perspectives: (1) Compared to the balanced diffusion Schrödinger bridge, our approach accurately recovers the correct growth and transition processes, eliminating false transitions caused by neglecting growth. (2) Compared to unbalanced optimal transport, the results obtained from the stochastic dynamics are more realistic, and we successfully learn the Waddington’s developmental landscape. We take $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$ in the following discussions.
6.1 Synthetic Dynamics
Inspired by (Sha et al., 2024), we adopt a same three-gene simulation model to explore the stochastic dynamics of gene regulation, as illustrated in Fig. 2(a). The system dynamics are governed by the following set of stochastic ordinary differential equations (ODEs):
Genes $X_{1}$ and $X_{2}$ mutually inhibit each other and self-activate, forming a toggle switch. An external signal $\beta$ uniformly activates both $X_{1}$ and $X_{2}$ independently of gene expression levels. Gene $X_{3}$ inhibits the expression of both $X_{1}$ and $X_{2}$. Here, $X_{i}(t)$ represents the concentration of gene $i$ at time $t$, with $\alpha_{i}$ and $\gamma_{i}$ denoting the strengths of self-activation and inhibition, respectively. The parameters $\delta_{i}$ represent gene degradation rates, while $\eta_{i}\xi_{t}$ accounts for the stochastic effects with additive white noise. The probability of cell division correlates positively with the expression of $X_{2}$, calculated as $g=\frac{X_{2}^{2}}{1+X_{2}^{2}}$. Upon division, cells inherit the gene expression states $(X_{1}(t),X_{2}(t),X_{3}(t))$ of the parent cell, subject to independent perturbations $\eta_{d}\mathcal{N}(0,1)$ for each gene, and transition independently thereafter.
The resulting gene regulatory and cellular dynamics are illustrated in Fig. 2, where a quiescent region and an area exhibiting both transition and growth can be observed. We focus on the projection of these dynamics onto the two-dimensional space of $(X_{1},X_{2})$ since on $X_{3}$ it remains quiescent.
(a) Illustration of gene regulatory network.
(b) The ground truth dynamics project on $(X_{1},X_{2})$.
Figure 2: Illustrative diagram of the gene regulatory network and the ground truth cellular dynamics project on $(X_{1},X_{2})$ coordinates.
(a) Balanced Schrödinger bridge (SF2M).
(b) Our DeepRUOT model.
Figure 3: Comparison of Our DeepRUOT with the balanced Schrödinger Bridge Method ($\sigma=0.25$).
In Fig. 3, we compare our proposed method with the balanced diffusion Schrödinger bridge method as described in (Tong et al., 2024b). Our findings indicate that neglecting the growth factor in the balanced diffusion Schrödinger bridge method leads to a false transition, incorrectly attracting cells in the quiescent state to the transition and growth region. The underlying reason for this discrepancy is that the growth factor causes a continuous increase in the number of cells. If the growth factor is ignored, the regions with increasing cell numbers will require more cells to transition into them to maintain balance, which ultimately leads to false transitions.
Our approach, which explicitly incorporates growth dynamics, successfully eliminates false transitions and achieves results consistent with the ground truth dynamics (Fig. 3(b)). Furthermore, the growth rates inferred by our method closely match the ground truth, demonstrating the robustness and accuracy of our unbalanced dynamic diffusion Schrödinger bridge model after normalization(Fig. 4).
(a) Ground truth growth rates.
(b) Growth rates learnt by our model.
Figure 4: Comparison of growth rates inferred by our DeepRUOT model with the ground truth growth rates.
6.2 Learning Waddington’s Developmental Landscapes
In both biology and physics, the Waddington’s landscape metaphor is a well-known model for representing cell fate decision process. This metaphor posits that meta-stable cell states correspond to wells in potential landscape, and state-transitions could be modelled as hopping between potential wells. Constructing such a potential landscape has been widely studied (Zhou and Li, 2016; Shi et al., 2022; Zhao et al., 2024; Li and Wang, 2013; Wang et al., 2010; Li and Wang, 2014; Bian et al., 2023, 2024; Zhou et al., 2024c, d), however it still remains a challenging problem using the single-cell omics data. The energy landscape is defined by
$U=-\sigma^{2}\log p_{\text{ss}}/2$
where $p_{\text{ss}}$ is the steady-state PDF satisfying the Fokker-Planck equation
$-\nabla\cdot(p_{ss}\boldsymbol{b})+\frac{\sigma^{2}}{2}\Delta p_{\text{ss}}=gp_{\text{ss}}.$
In this work, leveraging our model, we can naturally infer the time-varying potential energy landscape through the learned log density function, and use $-\sigma^{2}(t)\log p(\boldsymbol{x},t)/2$ represent the landscape at time $t$. The lower the energy function indicates more stable cell fates in landscape.
Figure 5: (a)The Waddington developmental landscape learnt at $t=1$ ($\sigma=0.25$). (b) The constructed landscape at $t=4$ ($\sigma=0.25$).
In Fig. 5, we can observe the temporal evolution of potential landscape in the synthetic gene regulation system. The quiescent cells always occupy a potential well, indicating its stability. Interestingly, the location of potential well belonging to transitional cell population are moving in the consistent direction with ground truth cell-state transition dynamics. Overall, the results suggest the accuracy and usefulness of DeepRUOT algorithm.
7 Conclusion
We have introduced a new deep learning method for learning regularized unbalanced optimal transport (RUOT) and continuous unbalanced stochastic dynamics from time-series snapshot data. By leveraging Fisher regularization, our method transforms a stochastic differential equation (SDE) problem into an ordinary differential equation (ODE) constraint. Through the use of neural network modeling for growth and death, our framework models dynamics without requiring prior knowledge of these processes. Additionally, it provides a new approach for learning RUOT problems.
Furthermore, we have summarized and explored the connections between RUOT and the Schrödinger bridge, specifically the branching Schrödinger bridge and the Schrödinger bridge on weighted paths, highlighting the associated challenges. We have demonstrated the effectiveness of our method on synthetic gene regulatory data, showing its ability to eliminate false transitions caused by neglecting growth processes.
Future directions involve extending the methodology to learn the cell latent embedding space jointly with the dynamics, and developing more computationally efficient algorithms. We expect future work to address these aspects and improve trajectory inference problems in various scenarios, including multi-omics data (e.g. single-cell ATAC-seq, spatial transcriptomics) and beyond biology in machine learning.
Acknowledgments
We thank Dr. Hugo Lavenant (Bocconi) for helpful discussions on Section 4.1 and Qiangwei Peng (PKU) for helpful suggestions on algorithm. We also thank Prof. Chunhe Li (Fudan), Juntan Liu (Fudan) and Prof. Qing Nie (UCI) for insightful discussions. We acknowledge the support from the High-performance Computing Platform of Peking University for computation.
Appendix A The equivalent formulation of RUOT with nonlinear Fokker-Planck equation constraints
In RUOT problem, when $\Psi\left({g}(\boldsymbol{x},t)\right)$ take the quadratic form i.e. $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$, one can define a formulation with nonlinear Fokker-Planck equation constraints as follows.
Definition A.1.
@@ Appendix A The equivalent formulation of RUOT with nonlinear Fokker-Planck equation constraints
To show its equivalence with RUOT, we study the property of its Fisher information regularisation form as below.
Theorem A.2.
When $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$, the regularized unbalanced optimal transport problem (3.9) is equivalent to
$\inf_{\left(p,\boldsymbol{v},\widetilde{g}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\alpha|\widetilde{g}(\boldsymbol{x},t)|^{2}-\frac{\sigma^{4}(t)}{16\alpha}(1+\log p)^{2}\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t$ (1.21)
where the infimum is taken all pairs $\left(p,\boldsymbol{v},\widetilde{g}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,T)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+\widetilde{g}(\boldsymbol{x},t)p+\frac{\sigma^{2}(t)}{4\alpha}(1+\log p)p.$ (1.22)
Proof A.3.
$\widetilde{g}=g-\frac{\sigma^{2}(t)}{4\alpha}(1+\log p).$ (1.23)
Substituting this variable into Eq. (3.9), then it follows immediately.
From Theorem A.2, we can find that if we introduce a new term in the continuity equation Eq. (3.8) as:
Then Eq. (1.22) now yields the standard continuity equation:
Appendix B Proof for Theorem 3.8
Proof B.1.
From Definition 3.7 we know that the Fokker-Planck equation for the SDE is

### v2

@@ D.3 RUOT with Nonlinear Fokker-Planck Equation Constraints
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+\widetilde{g}(\boldsymbol{x},t)p+\frac{\sigma^{2}(t)}{4\alpha}(1+\log p)p.$ (17)
coupled with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.
Proof.
$\widetilde{g}=g-\frac{\sigma^{2}(t)}{4\alpha}(1+\log p).$ (18)
Substituting this variable into Eq. (7), then it follows immediately. ∎
From Theorem D.1, we can find that if we introduce a new term in the continuity equation Eq. (6) as:
Then Eq. (17) now yields the standard continuity equation:
So we introduce a new definition.
Definition D.1.
@@
Remark D.1.
Note that when $\sigma(t)$ is constant Definition D.1 is consistent with the form provided in (Buze & Duong, 2023), and it has been conjectured that it is the higher order approximation scheme to the unbalanced optimal transport problem as $\sigma\rightarrow 0$.
D.4 RUOT in the General Case
For the general case, we consider the following equation
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)+gp,$ (19)
with the corresponding metric
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right)p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (20)
Definition D.2 (Regularized unbalanced optimal transport II).
Consider
$\inf_{\left(p,\boldsymbol{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right){p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (21)
where the infimum is taken all pairs $\left(p,\boldsymbol{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)+gp.$ (22)
coupled with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.
We similarly can reformulate Definition D.2 with the following Fisher information regularization.
Theorem D.2.
Consider the regularized unbalanced optimal transport problem Definition D.2, it is equivalent to
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\frac{1}{8}\left\|\boldsymbol{\sigma}(\boldsymbol{x},t)\nabla_{\boldsymbol{x}}\cdot\log\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ (23)
where the infimum is taken all pairs $\left(p,\boldsymbol{v},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+g(\boldsymbol{x},t)p.$ (24)
coupled with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.
Proof.
From Definition D.2 we know that the Fokker-Planck equation for the SDE is